# Field-free, Quasi-continuous Operation of Optical Nanofiber Interface with Two-dimensional Ferromagnetic Trap

### v3

@@ III Experimental demonstration
We take advantage of two-foil compactness (Fig. 1c) to construct a planar ONF-2D-MOT interface, schematically illustrated in Fig. 3. We further demonstrate transient absorption spectroscopy measurement with a measurement repetition rate of $f_{\rm rep}=250~{}$kHz. An anomalous line-broadening is found which is partly explained by the residual $B_{1}^{z}$ gradient intrinsic to the two-foil assembly (Fig. 2(a,b)).
@@ III.1 Experimental Setup
III.1.1 Two-foil vacuum assembly
The experimental setup is illustrated with Fig. 3a. A tapered nanofiber is fabricated with standard method  [32, 33, 34, 35]. The transmission of ONF is quite high, $T>85\%$. The fiber is epoxied to a titanium mount. The central $l_{0}\approx 5$ mm ONF section is centered to the mount, beneath which two pair of right-angle prisms are epoxied to the surface for the MOT-beam retro-reflections. The central pair of prisms are 101010~{}10mm high. The two side prisms are 666~{}6mm high. The ONF mounting assembly, carefully chosen to be non-magnetic, is connected to a CF35 stainless steel (SS-304) conflate flange. The ONF is optically feedthroughed to the vacuum with a Teflon interfacing technique [36]. The whole vacuum assembly is maintained to a $10^{-7}$ Pa base pressure. By opening a valve of a rubidium reservoir, a 87Rb partial pressure at $10^{-6}$ Pa level is achieved near ONF. To avoid Rb condensation, we gently heat the titanium assembly to $30^{\circ}$. Similarly, the ONF is heated by 200 $\rm{\bf\mu}W$ of 843843843~{}843nm light, (Fig. 4a) which is combined with the 780780780~{}780nm resonant probe beam with polarization optics and holographic gratings.
The two-foil assembly is constructed outside the vacuum with foil dimensions of $L=120~{}\text{mm}$ and $W=40~{}\text{mm}$. As shown in Fig. 5, each ${\bf\mu}-$metal foil is magnetized by a pair of cylindrical NdFeB magnets (25 mm diameter, 10 mm height) positioned 60 mm apart along the $z$ axis. An inter-foil gap of $2h=70~{}\text{mm}$ is used, which exceeds the ideal condition $h\ll W$ due to the vacuum tube diameter. The relatively thick foils, with $d=10~{}\text{mm}$, are advantageous for the 2D-MOT operation distance of $h=35~{}\text{mm}$ in this ex-vacuo design. To allow dynamic adjustment of the magnetic field, each ${\bf\mu}-$metal foil is wrapped with 23 turns of insulated copper wire with 1 mm diameter. The current passing through one foil is subsequently referred to as the “foil-coil” current (see Fig. 3d). Additionally, a coil with diameter $D=280~{}\text{mm}$ is mounted 30 mm away to provide a $B_{x}$ bias field, with its current denoted as the “x-coil” current.
We characterize the field distribution with a combination of numerical simulation (Fig. 2(a,b)) and Gauss-probe measurements (Fig. 5). As detailed in Appendix A, by adjusting in the simulation the the NdFeB strength, the numerical results can match quite well with the measurements. The surface field near $S_{2}$ is highly uniform and stable (Fig. 5). For the two-foil assembly, we find a quasi-2D magnetic field-zero line near the ONF location with field gradients of $B_{1}^{x}\approx 21~{}{\rm G/cm}$, $B_{1}^{y}\approx-24~{}{\rm G/cm}$, $B_{1}^{z}\approx 3~{}{\rm G/cm}$ along $x,y,z$ respectively.
@@ III.1.2 Prism MOT
[…]
@@ III.2 MOT loading
The MOT loading process is monitored by the ONF fluorescence. As illustrated with Fig. 3d setup, we direct an ONF output to a holographic grating to select the 780 nm fluorescence for single-photon counting (SPCM-AQRH, Excelitas). We switch off and back on the MOT beams with an acousto-optical modulator (AOM) to record the MOT-filling process. From the camera image, the final length of 2D-MOT along $z$ and in the $x-y$ plane are $l=4~{}$mm and $\Delta x=0.8~{}$mm respectively.
The fluorescence counting rate curve as shown in Fig. 3e includes a rapid initial growth which is followed by a slow increase further. The underlying mechanism is explained as following: Magneto-optical trap is known to have an irregular density distribution [18] associated with the quite unavoidable variations of intensity-imbalance between counter-propagating MOT beams. During the initial MOT loading, atoms prefer to be accumulated in intensity-balanced locations. Later, when there are plenty of atoms in these locations, multi-scattering starts to drive their expansions, leading to a large, “density-limited ” sample [20] with a uniform local density $\rho_{0}$. Here, the rapid initially growth of ONF fluorescence is associated with filling of intensity-balanced spots across ONF. The slow increase of ONF fluorescence later is associated with the multi-scattering-driven spreading to uniformly fill the ONF interaction zone, which effectively increase the interaction length $l$. We fit the fluorescence curve with a phenomenological model to include this multi-scattering-driven MOT expansion dynamics, detailed in Appendix B. The multi-scattering-limited density $\rho_{0}$ is derived from ONF absorption measurement. In particular, we find that the fully loaded atomic sample attenuates the ONF-guided resonant probe with an ${\rm OD}=2.2$ optical depth (see e.g. Fig. 4b), which suggests a uniform $\rho_{0}=4\times 10^{10}$/cm3 over the $l=4~{}$mm interaction length. In Fig. 3e we also estimate $N_{\rm eff}={\rm OD}(t)/{\rm od}_{1}$ as an effective number of atoms coupled to ONF. Here ${\rm OD}(t)$ is scaled with the fluorescence counting rate during the MOT loading. The ${\rm od}_{1}=0.09$ is single-atom resonant optical depth at a 100100100~{}100nm distance away from the $d=500~{}$nm ONF [2, 37].
@@ III.3 Field control
We shift the 2D-MOT along $x$ and $y$ by adjusting the $x$-coil and foil-coil currents, demonstrating this field control by measuring the ONF-coupled steady-state MOT fluorescence, as shown in Fig. 3f.
We begin by optimizing the ONF fluorescence counting rate through adjustments of both the $x$-coil and foil-coil currents. With the foil-coil current then fixed at its optimal 3.5 A, we scan the $x$-coil current in small steps between 0.76 A and 1.5 A. This procedure shifts the 2D-MOT along $x$ by approximately 1 mm, as indicated by the mirror images observed on the camera (upper-most and lower-most MOT images in Fig. 3c). Since the 2D-MOT has a transverse size of about 0.8 mm, this shift is sufficient to toggle the ONF-coupled fluorescence off and back on, as shown by the red markers in Fig. 3f.
Next, with the $x$-coil current set to its optimal 1.16 A, we scan the foil-coil current between 1.2 A and 5.4 A. This adjustment shifts both the 2D-MOT and its second-order image transversely by approximately 1.8 mm on the camera (the two central images in Fig. 3c). The resulting variations in the ONF fluorescence counting rate are represented by the blue markers in Fig. 3f.
From the data in Fig. 3f, we estimate the foil-coil current tuning response to be approximately 1 G/A at the 2D-MOT location. Notably, the field is enhanced by a factor of $\eta\approx 5$ due to the magnetization of the ${\bf\mu}$-metal. Without the ${\bf\mu}$-metal, the same foil-coil would produce only about 0.2 Gauss per Amp at the same distance. This soft-magnetic field enhancement leads to the low power operation. In this setup, $\eta\approx 5$ is constrained by the moderate $d/W$ ratio and the resultant strong depolarization field [25]. In future work, a shorter working distance $h$ and a thinner foil with a smaller $d/W$ ratio could reduce demagnetization, potentially increasing $\eta$ to the hundred level. Regardless of the $\eta$ value, the soft-magnetization protects field uniformity against the foil-coil winding imperfection, as being explained in Sec. II.1.
@@ III.4 High-rep-rate transient ONF spectroscopy
Figure 4: (a): Absorption spectroscopy setup and timing sequence. SPCM: Single photon counting module. Here $\tau_{\rm c}=30~{}$ns. $\tau_{\rm p}=200~{}$ns, 180 ns in (b) and (c,d) respectively. (b): Typical SPCM readout of the ONF-transmitted probe (red) and reference (blue). (c): Steady-state absorption spectroscopy and Lorentzian fit. (d): Transmission of a resonant probe ($\Delta=0$) from (b) with fit according to linear dispersion theory  [38, 39, 25, 40]. The dashed box suggests the range of data for the average to compile the $\Delta=0$ data point in (c).
As illustrated in Fig. 2(a,b), even with a working distance of $h=35$ mm, the field distribution in the two-foil configuration remains highly elongated along $z$. The 3 Gauss/cm gradient leads to 0.6 Gauss field strength deviation across the $l=4$ mm ONF interaction section aligned to the quasi-zero-field line. This level of field suppression should be sufficient for atomic spectroscopy at moderate precision, allowing measurements without globally switching off the field.
@@
During the probe window, we generate a composite probe pulse using a fiber electro-optical modulator (fEOM)-based optical arbitrary waveform generator (OAWG) [41]. The pulse includes a 30 ns “calibration pulse” followed by a $\sim 200$ ns “interrogation pulse”. The calibration pulse is 400 MHz blue-detuned from the $F=2\rightarrow F^{\prime}=3$ hyperfine transition, leaving its ONF transmission unaffected by the presence of cold atoms. The interrogation pulse is scanned in detuning $\Delta$ from -20 MHz to 20 MHz. We keep the probe power at a low level of 5 pW to avoid saturating the atomic transition [42]. For each detuning $\Delta$, we integrate the single-photon counting module (SPCM) measurements over $T_{i}=10$ seconds. The resulting single-photon counts from all $f_{\rm rep}\times T_{i}=2.5\times 10^{6}$ measurements are histogrammed to yield the ONF-transmitted probe signal $I_{p}(t)$, shown in red in Fig. 4b. We note the MOT fluorescence is visible in $I_{p}(t)$ beyond the probe time window, as expected.
@@
III.5 Anomalous line broadening
With the transient transmission data $T(t)$ as those in Fig. 4d, we retrieve the near-field coupled atomic response in two ways. First, as suggested by the green dashed box, we take the $T(t)$ data between $175~{}{\rm ns}<t<180~{}{\rm ns}$ and calculate the average $\bar{T}$ at the steady state. We then evaluate the optical depth of the evanescently coupled cold atoms according to Beer-Lambert law [43], ${\rm OD}=-{\rm ln}(\bar{T})$, at each probe detuning $\Delta$. Typical absorption spectroscopy data are given in Fig. 4c together with Lorentzian fit. In the second method, we choose a particular $T(\Delta,t)$ curve as those in Fig. 4b, and apply linear dispersion theory [38, 39, 25, 40] to fit the Lorentzian atomic response [44]. In both cases, we arrive at a fitted linewidth of $\tilde{\Gamma}\sim 2\pi\times 9\sim 10~{}$MHz. This value is substantially larger than one might expect from the $\Gamma=2\pi\times 6.1~{}$MHz natural linewidth of the D2 transition [45], even after accounting for the ONF-enhanced emission and the surface interactions [11, 17, 8].
We checked a list of potential systematics that may lead to the anomalous line broadening. First, it is worth noting that unlike previous spectroscopic work where a prolonged measurement could affect the atomic trajectories [10, 17, 8], here, the atomic motion according to MOT dynamics is effectively frozen for our nanosecond probe. By increasing the probe power and repeating the measurements, we verified that our probe pulse is well below saturation in the near field [42]. To verify that cooling is effective between the nanosecond probes, the results were also repeated by increasing $T_{\rm rep}$ to 8 ${\bf\mu}$s and 16 ${\bf\mu}$s. Finally, the linewidth is not noticeably affected by long-range resonant dipole interaction [3], which is verified by repeating the observation with reduced atom numbers.
The observed linewidth broadening is partly explained by residual magnetic field. In particular, we find numerically [46, 37] that coherent coupling of different Zeeman sublevels by Gauss-level field, perpendicular to the quantization axis set by light, leads to notable line-broadening. However, by shifting the 2D-MOT center in the $x-y$ plane as in Fig. 3f, and by scanning a $z-$bias field, we verified that the $\tilde{\Gamma}=2\pi\times 9\sim 10~{}$MHz linewidth is indeed as small as we can get. We also note the 2D-MOT and the nanofiber are parallel to each other within a few degrees (Fig. 3c). With ONF aligned to the field-zero line, the $B_{1}^{z}\approx 3~{}$G/cm gradient should merely lead to $\delta B=0.6$ G deviation of field strength. The average field strength is therefore at 0.3 G level, much less than what is required to broaden the line even by just 1 MHz, about 1 G, according to numerical simulations [46, 37].
Therefore, even though the $\tilde{\Gamma}\approx 2\pi\times 9\sim 10$ MHz linewidth could be partly explained by the residual magnetic field in our setup, it remains a puzzle to be fully explained. A detailed account of our investigation on this topic will be given in a future publication [47].

### v5

@@ III Experimental demonstration
We take advantage of two-plate compactness (Fig. 1c) to construct a planar ONF-2D-MOT interface, schematically illustrated in Fig. 3. We further demonstrate transient absorption spectroscopy measurement with a measurement repetition rate of $f_{\rm rep}=250$~kHz.
@@ III.1 Experimental Setup
III.1.1 Two-plate vacuum assembly
The experimental setup is illustrated with Fig. 3a. A tapered nanofiber is fabricated with standard method  [34, 35, 36, 37]. The transmission of ONF is quite high, $T>85\%$. The fiber is epoxied to a titanium mount. The central $l_{0}\approx 5$ mm ONF section is centered to the mount, beneath which two pair of right-angle prisms are epoxied to the surface for the MOT-beam retro-reflections. The central pair of prisms are 1010~$m$ high. The two side prisms are 1010~$m$ high too. The ONF mounting assembly, carefully chosen to be non-magnetic, is connected to a CF35 stainless steel (SS-304) conflate flange. The ONF is optically feedthroughed to the vacuum with a Teflon interfacing technique [38]. The whole vacuum assembly is maintained to a $10^{-7}$ Pa base pressure. By opening a valve of a rubidium reservoir, a 87Rb partial pressure at $10^{-6}$ Pa level is achieved near ONF. To avoid Rb condensation, we gently heat the titanium assembly to $30^{\circ}$C. Similarly, the ONF is heated by 100 $\rm{\bf\mu}W$ of 843843~nm light, (Fig. 4a) which is combined with the 780780~nm resonant probe beam with polarization optics and holographic gratings.
The two-plate assembly is constructed outside the vacuum with plate dimensions of $L=120~\text{mm}$ and $W=40~\text{mm}$. As shown in Fig. 5, each $\bf\mu$}-metal plate is magnetized by a pair of cylindrical NdFeB magnets (25 mm diameter, 10 mm height) positioned 60 mm apart along the $z$ axis. An inter-plate gap of $2h=70~\text{mm}$ is used, which exceeds the ideal condition $h\ll W$ due to the vacuum tube diameter. The relatively thick plates, with $d=10~\text{mm}$, are advantageous for the 2D-MOT operation distance of $h=35~\text{mm}$ in this ex-vacuo design. To allow dynamic adjustment of the magnetic field, each $\bf\mu$}-metal plate is wrapped with 23 turns of insulated copper wire with 1 mm diameter. The current passing through one plate is subsequently referred to as the “plate-coil” current (see Fig. 3d). Additionally, a coil with diameter $D=280~\text{mm}$ is mounted 35 mm away to provide a $B_{x}$ bias field, with its current denoted as the “x-coil” current. Simulation with Radia [29, 30] (Appendix A) confirms that the offset field generated by the $x$-coil is uniform across ONF to within $1\%$.
We characterize the field distribution with a combination of numerical simulation (Fig. 2(a,b)) and Gauss-probe measurements (Fig. 5). As detailed in Appendix A, by adjusting in the simulation the the NdFeB strength, the numerical results can match quite well with the measurements. The surface field near $S_{2}$ is highly uniform and stable (Fig. 5). For the two-plate assembly, we find a quasi-2D magnetic field-zero line near the ONF location with field gradients of $B_{1}^{x}\approx 21~{\rm G/cm}$, $B_{1}^{y}\approx-24~{\rm G/cm}$, $B_{1}^{z}\approx 3~{\rm G/cm}$ along $x,y,z$ respectively.
@@ III.1.2 Prism MOT
As shown in Fig. 3a, a MOT beam of $18~\text{mm}$ diameter is directed through the vacuum viewport toward the central right-angle prism pair atop the titanium assembly. The laser beam, with power $P=20~\text{mW}$, is detuned by $\Delta=-2\Gamma$ ($12~\text{MHz}$) from the $F=2\rightarrow F^{\prime}=3$ D2 hyperfine transition of 87Rb. It is circularly polarized to align with the quadrupole field direction, as depicted in Fig. 1c. Additionally, $1~\text{mW}$ of laser power resonant with the $F=1\rightarrow F^{\prime}=2$ hyperfine transition is combined with the MOT beam to provide hyperfine repumping. To prevent the ONF from encountering a dark line in the retro-reflected beam, caused by a small gap between the prisms, we slightly tilt the beam’s incidence angle.
To provide axial cooling and trapping, two additional $z$-beams are sent through the same viewport and redirected by the pair of smaller prisms to intersect at the ONF. These $z$-beams have a smaller diameter of $3~\text{mm}$, each with a power of $P=3~\text{mW}$. With a weak axial magnetic gradient $B_{1}^{z}\approx 3~\text{G/cm}$, we find the polarizations of the $z$-beams to be noncritical for 2D-MOT operation. Setting the $z$-beam circular polarizations in line with the field direction slightly compresses the MOT along the $z$-axis, as expected. However, in our setup, we intentionally choose the polarizations as opposite to the field direction, which helps to extend the 2D-MOT length by approximately $20\%$. A typical MOT image is shown in Fig. 3c, together with three mirror images.
@@ III.2 MOT loading
The MOT loading process is monitored by the ONF fluorescence. As illustrated with Fig. 3d setup, we direct an ONF output to a holographic grating to select the 780 nm fluorescence for single-photon counting (SPCM-AQRH, Excelitas). We switch off and back on the MOT beams with an acousto-optical modulator (AOM) to record the MOT-filling process. From the camera image, the final length of 2D-MOT along $z$ and in the xyx-y plane are l=4l=4~$m$ and $\Delta x=0.8$~$m$ respectively.
The fluorescence counting rate curve as shown in Fig. 3e includes a rapid initial growth which is followed by a slow increase further. The underlying mechanism is explained as following: Magneto-optical trap is known to have an irregular density distribution [18] associated with the quite unavoidable variations of intensity-imbalance between counter-propagating MOT beams. During the initial MOT loading, atoms prefer to be accumulated in intensity-balanced locations. Later, when there are plenty of atoms in these locations, multi-scattering starts to drive their expansions, leading to a large, “density-limited ” sample [20] with a uniform local density $\rho_{0}$. Here, the rapid initially growth of ONF fluorescence is associated with filling of intensity-balanced spots across ONF. The slow increase of ONF fluorescence later is associated with the multi-scattering-driven spreading to uniformly fill the ONF interaction zone, which effectively increase the interaction length $l$. We fit the fluorescence curve with a phenomenological model to include this multi-scattering-driven MOT expansion dynamics, detailed in Appendix B. The multi-scattering-limited density $\rho_{0}$ is derived from ONF absorption measurement. In particular, we find that the fully loaded atomic sample attenuates the ONF-guided resonant probe with an ${\rm OD}=2.2$ optical depth (see e.g. Fig. 4b), which suggests a uniform $\rho_{0}=3\times 10^{10}$/cm3 over the l=4l=4~$m$ interaction length. In Fig. 3e we also estimate $N_{\rm eff}={\rm OD}(t)/{\rm od}_{1}$ as an effective number of atoms coupled to ONF. Here ${\rm OD}(t)$ is scaled with the fluorescence counting rate during the MOT loading. The ${\rm od}_{1}=0.09$ is single-atom resonant optical depth at a 125125~nm distance away from the d=500d=500~nm ONF [2, 39] in our setup (see Appendix B).
@@ III.3 Field control
We shift the 2D-MOT along $x$ and $y$ by adjusting the $x$-coil and plate-coil currents, demonstrating this field control by measuring the ONF-coupled steady-state MOT fluorescence, as shown in Fig. 3f.
We begin by optimizing the ONF fluorescence counting rate through adjustments of both the $x$-coil and plate-coil currents. With the plate-coil current then fixed at its optimal 3.5 A, we scan the $x$-coil current in small steps between 0.76 A and 1.5 A. This procedure shifts the 2D-MOT along $x$ by approximately 1 mm, as indicated by the mirror images observed on the camera (upper-most and lower-most MOT images in Fig. 3c). Since the 2D-MOT has a transverse size of about 0.8 mm, this shift is sufficient to toggle the ONF-coupled fluorescence off and back on, as shown by the green markers in Fig. 3f.
Next, with the $x$-coil current set to its optimal 1.16 A, we scan the plate-coil current between 1.2 A and 5.4 A. This adjustment shifts both the 2D-MOT and its second-order image transversely by approximately 1.45 mm on the camera (the two central images in Fig. 3c). The resulting variations in the ONF fluorescence counting rate are represented by the red markers in Fig. 3f.
From the data in Fig. 3f, we estimate the plate-coil current tuning response to be approximately 0.8 G/A at the 2D-MOT location. Notably, the field is enhanced by a factor of $\eta\approx 4$ due to the magnetization of the $\bf\mu$}-metal. Without the $\bf\mu$}-metal, the same plate-coil would produce only about 0.2 Gauss per Amp at the same distance. This soft-magnetic field enhancement leads to the low power operation. In this setup, $\eta\approx 4$ is constrained by the moderate $d/W$ ratio and the resultant strong depolarization field [25]. In future work, a shorter working distance $h$ and a thinner plate with a smaller $d/W$ ratio could reduce demagnetization, potentially increasing $\eta$ to the hundred level. Regardless of the $\eta$ value, the soft-magnetization protects field uniformity against the plate-coil winding imperfection, as being explained in Sec. II.1.
@@ III.4 High-rep-rate transient ONF spectroscopy
Figure 4: (a): Absorption spectroscopy setup and timing sequence. SPCM: Single photon counting module. Here $\tau_{\rm c}=30$~ns. $\tau_{\rm p}=200$~ns in (b) and (c,d) respectively. (b): Typical SPCM readout of the ONF-transmitted probe (red) and reference (blue). (c): Steady-state absorption spectroscopy and Lorentzian fit. (d): Transmission of a resonant probe ($\Delta=0$) from (b) with fit according to linear dispersion theory  [40, 41, 25, 42] (Appendix C). The dashed box suggests the range of data for the average to compile the $\Delta=0$ data point in (c).
As illustrated in Fig. 2(a,b), even with a working distance of $h=35$ mm, the field distribution in the two-plate configuration remains highly elongated along $z$. The 3 Gauss/cm gradient leads to 0.6 Gauss field strength deviation across the $l=4$ mm ONF interaction section aligned to the quasi-zero-field line. This level of field suppression should be sufficient for atomic spectroscopy at moderate precision, allowing measurements without globally switching off the field.
@@
During the probe window, we generate a composite probe pulse using a fiber electro-optical modulator (fEOM)-based optical arbitrary waveform generator (OAWG) [43]. The pulse includes a 30 ns “calibration pulse” followed by a $\sim 200$ ns “interrogation pulse”. The calibration pulse is 400 MHz blue-detuned from the $F=2\rightarrow F^{\prime}=3$ hyperfine transition, leaving its ONF transmission unaffected by the presence of cold atoms. The interrogation pulse is scanned in detuning $\Delta$ from -20 MHz to 20 MHz. We keep the probe power at a low level of 5 pW to avoid saturating the atomic transition [44]. For each detuning $\Delta$, we integrate the single-photon counting module (SPCM) measurements over $T_{i}=60$ seconds. The resulting single-photon counts from all $f_{\rm rep}\times T_{i}=1.5\times 10^{7}$ measurements are histogrammed to yield the ONF-transmitted probe signal $I_{p}(t)$, shown in red in Fig. 4b. We note the MOT fluorescence is visible in $I_{p}(t)$ beyond the probe time window, as expected.
@@
III.5 Linewidth measurement
With the transient transmission data $T(t)$ as those in Fig. 4d, we retrieve the near-field coupled atomic response in two ways (Appendix C). First, as suggested by the green dashed box, we take the $T(t)$ data between $195~{\rm ns}<t<200~{\rm ns}$ and calculate the average $\bar{T}$ at the steady state. We then evaluate the optical depth of the evanescently coupled cold atoms according to Beer-Lambert law [45], ${\rm OD}=-{\rm ln}(\bar{T})$, at each probe detuning $\Delta$. Typical absorption spectroscopy data are given in Fig. 4c together with Lorentzian fit. In the second method, we choose a particular $T(\Delta,t)$ curve as those in Fig. 4b, and apply linear dispersion theory [40, 41, 25, 42, 46] to fit the Lorentzian atomic response. In both cases, we arrive at a fitted linewidth of $\tilde{\Gamma}/2\pi\approx 9.5\sim 10.5$~MHz. This value is substantially larger than one might expect from the $\Gamma=2\pi\times 6.1$~MHz natural linewidth of the D2 transition [47], even after accounting for the ONF-enhanced emission and the surface interactions [11, 17, 8].
It is worth noting that unlike previous spectroscopic work with free atoms [15, 17, 8], where a prolonged measurement could affect the atomic trajectories near the nanofiber, here the atomic motion is effectively frozen for our nanosecond probe. As we expect a largely uniform atomic density in the near field, the spectroscopic measurement can be quite conveniently interpreted to infer surface interactions [48]. Therefore, that we see a Lorenzian absorption profile as in Fig. 4(c,d), broadened but without substantial van-der-Waals-shift induced asymmetry [15, 49, 17, 8], is quite puzzling. In Appendix C we detail the Fig. 4(c,d) measurements, and provide strong evidence that the anomalously large linewidth as in Fig. 4(c,d) is affected by the surface quality of ONF [49] and is likely of magnetic origin.
Here, we note that the observed linewidth broadening can hardly be attributed to the residual field of our two-plate assembly alone. In particular, when the ONF is aligned along the field-zero line, the measured gradient of $B_{1}^{z}\approx 3$ G/cm should lead to a field variation of only $\delta B=0.6$ G. The resulting average field strength would therefore remain at the 0.3 G level, far below the required field to broaden the line by even 1 MHz (approximately 1 G), according to numerical simulations [50, 39] (Appendix C.2). Instead, explaining the observed broadening would require $B_{1}^{z}\approx 15$ G/cm, which is clearly inconsistent with the 2D-MOT observations in Sec. II.5.